OpenStudy (anonymous):

does the sequence n*tan(1/n) converge or diverge? @OCW Scholar - Sin…

OpenStudy (anonymous):

limit test, series diverges.

OpenStudy (jamesj):

Guys, the question is does the limit as \( n \rightarrow \infty\) of n tan(1/n) exist. Note first that \[ n \tan(1/n) = \frac{\sin(1/n)}{1/n} \sec(1/n) \] Now this limit does exist. The limit of the first term is 1 and the limit of the sec term is also 1.

OpenStudy (anonymous):

You don't need to work so hard. Since we're looking at n*tan(1/n) as n --> ∞ : Assume x = 1/n That means we're looking for tan(x)/x as x approaches 0. Using L'Hopitals rule, the solution is:\[\sec^2(x)/1 = 1\]

OpenStudy (jamesj):

a) Both our proofs are just a couple of lines b) One of uses artillery while the other just a sling shot. I think it's clear which is more elegant :-)

OpenStudy (anonymous):

as james said the given question is \[\lim_{n \rightarrow \infty} n \tan (1/n)\] read it as \[\lim_{n \rightarrow \infty} \tan(1/n)/(1/n)\] lets make a substitution to solve the problems easier make the substitution \[n \rightarrow 1/x \] as \[n \rightarrow \infty \implies x \rightarrow 0\] so the limit changes to \[\lim_{x \rightarrow 0} \tan x/x\] which is 1